A local country club has a membership of 600 and operates facilities that include an 18-hole championship golf course and 12 tennis courts. Before deciding whether to accept new members, the club president would like to know how many members regularly use each facility. A survey of the membership indicates that 65% regularly use the golf course, 49% regularly use the tennis courts, and 7% use neither of these facilities regularly. What is the probability that a worker regularly uses both golf and tennis facilities?
A. .49
B. .21
C. .70
D. .93

Found 3 solutions by edjones, pharoah99, richwmiller:
Answer by edjones(8007)

Let uses both = x
Only golf = .65-x
only tennis= .49-x
uses neither= .07
.
x+.65-x+.49-x+.07=1.00
1.21-x=1.00
-x=-.21
x=.21
.
Ed

Answer by pharoah99(1)
Answer by richwmiller(17219)

You jumped from talking about members to asking about workers!
We know nothing about worker privileges nor habits.